OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 64. What is the length of the altitude?

OpenStudy (anonymous):

@jim_thompson5910 hellppp

jimthompson5910 (jim_thompson5910):

One sec while I draw it

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

@jim_thompson5910 you almost done?

jimthompson5910 (jim_thompson5910):

yes one moment

OpenStudy (anonymous):

ok:)

jhonyy9 (jhonyy9):

HELP001 do you know the thesis of altitude in case of a right triangle ?

jimthompson5910 (jim_thompson5910):

If you draw out the problem, you'll get something like this

OpenStudy (anonymous):

no @jhonyy9

jhonyy9 (jhonyy9):

you dont have learned it in your schoole ?

jimthompson5910 (jim_thompson5910):

So we basically have 2 smaller right triangles (along with a large right triangle) The smaller right triangle on the left side has legs of h and 16 units. It also has a hypotenuse of y. So we can say h^2 + 16^2 = y^2 ------------------------------------------------------- The smaller right triangle on the right side has legs of h and 64 units. It also has a hypotenuse of x. So we can also say h^2 + 64^2 = x^2

jimthompson5910 (jim_thompson5910):

Finally, the larger right triangle has legs of x and y along with a hypotenuse of 64+16 = 80 So we can say x^2 + y^2 = 80^2

jimthompson5910 (jim_thompson5910):

So we now have 3 non-linear equations with 3 unknowns h^2 + 16^2 = y^2 h^2 + 64^2 = x^2 x^2 + y^2 = 80^2

OpenStudy (anonymous):

so the anwser is 16? :) @jim_thompson5910

OpenStudy (anonymous):

8 4 32 16 theses are the possible anwsers!

jimthompson5910 (jim_thompson5910):

no, let's find solve for h x^2 + y^2 = 80^2 (h^2 + 64^2 ) + y^2 = 80^2 (h^2 + 64^2 ) + (h^2 + 16^2) = 80^2 2h^2 + 64^2+16^2 = 80^2 2h^2 + 4096 + 256 = 6400 I'll let you finish and solve for h

jimthompson5910 (jim_thompson5910):

Let me know if you get stuck

jhonyy9 (jhonyy9):

@jim_thompson5910 you dont like explain inthis case the thesis of altitude inside a right triangle --- please ! i think is more easy understanding how we need solveing this exersice

OpenStudy (anonymous):

i got 32 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it HELP001, nice work

jimthompson5910 (jim_thompson5910):

jhonyy9 I'm not sure what you're saying

jhonyy9 (jhonyy9):

ok look please the thesis of altitude inside a right triangle say that the length squared of altitude from the right angle on hypothenuse is equal with the product of this two segments of hypothenuse

jhonyy9 (jhonyy9):

like in this case h^2 =16*64 h^2=4^2 *8^2 h=4*8 h=32 i think this is more easy in this way

jimthompson5910 (jim_thompson5910):

maybe but you have to know the formula to get that

jimthompson5910 (jim_thompson5910):

so i guess it depends

jhonyy9 (jhonyy9):

yes is right but this is a thesis what i have learned in my schoole

jhonyy9 (jhonyy9):

HELP001 how you think it ,you dont learned again ?